# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ Abstract
In this paper, we investigate the phenomenon of coming down from infinity for (sub)critical cooperative branching processes with pairwise interactions (BPI processes for short) under appropriate conditions. BPI processes are continuous-time Markov chains that extend pure branching dynamics by incorporating additional mechanisms that allow both competition and cooperation events between pairs of individuals.
Specifically, we focus on characterising the speed at which BPI processes evolve when starting from a very large initial population in the subcritical and critical cooperative regimes. Further, in the subcritical cooperative regime, we analyse their second-order fluctuations.
@@ 1 Introduction
In this paper, we explore the speed of coming down from infinity for both subcritical and critical cooperative branching processes with pairwise interactions (referred to as BPI processes), under appropriate conditions. These processes were recently introduced in their current form by González-Casanova et al. [14] and further analysed by the authors in [7]. Our primary focus is on understanding the short-time behaviour of BPI processes when the initial population size is significantly large.
BPI processes emerge as a natural extension of the family of probabilistic models known as branching processes, addressing some of their inherent limitations, particularly their unrealistic assumptions. By incorporating pairwise interactions such as competition and cooperation between individuals, BPI processes can capture complex dynamics observed in biological systems. This generalisation of BGW processes has been studied in various contexts by several authors, including Jagers [20], Kalinkin [21, 22, 23, 24, 25], Lambert [28], and more recently by González-Casanova et al. [14] and the authors [7]. In these works, the birth and death rates are generalised by considering polynomial rates as functions of the population size which can be interpreted as different types of interactions between individuals.
Formally, let $\mathbb{N}\coloneqq\{1,2,\ldots\}$ and $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$. BPI processes are continuous-time Markov chains with state space $\mathbb{N}_{0}$. They feature natural births and deaths at rates proportional to the current population size, similar to the continuous-time Bienaymé-Galton-Watson (BGW for short) process, but also incorporate additional birth and death events due to cooperation and competition, occurring at rates proportional to the square of the population size. Specifically, a BPI process with parameters $c,d\in\mathbb{R}_{+}:=[0,\infty)$ and sequences $(b_{i})_{i\geq 1}$ and $(\pi_{i})_{i\geq 1}$ where $b_{i},\pi_{i}\in\mathbb{R}_{+}$, for $i\geq 1$, and $b\coloneqq\sum_{i\geq 1}b_{i}<\infty$ and $\rho\coloneqq\sum_{i\geq 1}\pi_{i}<\infty$, is defined by the infinitesimal generator $Q=(q_{i,j})_{i,j\in\mathbb{N}_{0}}$ given by
$q_{i,j}=\left\{\begin{array}[]{ll}i\pi_{j-i}+i(i-1)b_{j-i}&\mbox{ if }i\geq 1\,\,\text{and}\,\,j>i,\\ di+ci(i-1)&\mbox{ if }i\geq 1\,\,\text{and}\,\,j=i-1,\\ -i(d+\rho+(b+c)(i-1))&\mbox{ if }i\geq 1\,\,\text{and}\,\,j=i,\\ 0&\text{otherwise}.\\ \end{array}\right.$ (5)
@@
From the infinitesimal generator (5), or $Q$-matrix, we observe that the additional death rate $cz(z-1)$ represents competition, where one of the $z$ individuals selects another from the remaining $z-1$ at a constant rate $c$ and kills it. The additional birth rate $bz(z-1)$ reflects cooperative interactions, where pairs of individuals produce $i$ new individuals with probability $b_{i}/b$.
@@
According to González-Casanova et al. [14], this quantity significantly influences the long-term behaviour of BPI processes. A BPI process is termed supercritical, critical, or subcritical cooperative depending on whether $\mathbf{m}_{\rm int}>0$, $\mathbf{m}_{\rm int}=0$ or $\mathbf{m}_{\rm int}<0$, respectively. Theorem 1 in [14] shows that in the supercritical regime, and when
$\mathbf{m}_{\rm br}\coloneqq-d+\sum_{i\geq 1}i\pi_{i}<\infty,$
@@
the associated BPI process may explode in finite time with positive probability. In the subcritical and critical cooperative regime, and when $\mathbf{m}_{\rm br}<\infty$, the associated BPI process does not explode in finite time a.s. More recently, the authors [7] studied the explosion event for general subcritical and critical cooperative BPI processes (i.e. not necessarily satisfying that $\mathbf{m}_{\rm br}<\infty$) and provide general integral test, see Theorems 1 and 2 in [7] for further details.
@@
When $c=b=0$, the BPI process reduces to the BGW process. A special case of BPI processes is the so-called logistic branching process (LB-process), deeply studied in [28], which corresponds to $c>0$ and $b=0$. Another noteworthy example is the counting block process of the renowned Kingman coalescent which arises when the BPI process does not consider natural deaths, branching, and cooperation events (i.e., $c>0$ and $d=\rho=b=0$); see [26]. More generally, when $c,\rho>0$ and $d=b=0$, the BPI-process is exactly the block counting process of an exchangeable fragmentation-coagulation (EFC for short) process, introduced in [5], where the coagulation mechanism features only the Kingman component. Moreover, the BPI process can be viewed as the counting process of a particle system where particles can coagulate (via competition interactions), fragment without collision (via branching), or fragment due to collisions (via cooperative interactions). Ultimately, when $\pi_{i}=b_{i}=0$ for all $i\geq 2$, the BPI process reduces to a classic birth-and-death process.
@@
Motivated by studies on population dynamics with initially large populations, the phenomenon of coming down from infinity has garnered significant attention. This phenomenon describes a scenario where a random process starting from an infinite number of individuals (or blocks) immediately reaches a finite state. Kingman [26] was the first to observe this phenomenon in the context of the Kingman coalescent. Subsequently, Schweinsberg [34] provided necessary and sufficient conditions for the occurrence of this phenomenon in the setting of the $\Lambda$-coalescent. Further work, for the $\Lambda$-coalescent, includes analyses of the speed of coming down from infinity by Berestycki et al. [6] and second-order asymptotics by Limic and Talarczyk [29, 30]. The case of EFC processes was treated in Berestycki [5] where a sufficient condition for the process to come down from infinity is provided. Recent studies include those by Foucart [12], exploring the phenomenon of coming down from infinity for EFC processes with multiple coagulations but not simultaneous, as in the $\Lambda$-coalescent, and the fragmentation dislocates, at a finite rate, an individual block into sub-blocks of infinite size. Bansaye et al. [4] described the speed of coming-down from infinity for birth-and-death processes that eventually become extinct. More recently, Bansaye [3] developed a general framework for the phenomenon across various stochastic processes with jumps described by stochastic differential equations. Finally, it is important to note that the phenomenon of coming down from infinity for subcritical cooperative BPI processes was first studied by González-Casanova et al. [14] under the assumption that $\mathbf{m}_{\rm br}<\infty.$ More recently, the authors [7] established a sufficient condition for general subcritical and critical cooperative BPI processes to come down from infinity (see Section 2.3). However, nothing is known about the speed at which BPI processes come down from infinity.
@@
Our aim is to determine the speed at which a subcritical or critical cooperative BPI process comes down from infinity. In the subcritical cooperative regime (i.e. $\mathbf{m}_{\rm int}<0$), under the assumption that $\mathbf{m}_{\rm br}<\infty$ and the cooperation term $(b_{i})_{i\geq 1}$ has positive moments in $(1,2]$ we show that the rate is asymptotically equivalent to $-\mathbf{m}_{\rm int}t^{-1}$, similarly as in the Kingman coalescent case where $-\mathbf{m}_{\rm int}=c$. In particular, we have that the speed of coming down from infinity of the block counting process of the EFC process with only Kingman component is asymptotically equivalent to $ct^{-1}$.
In the critical cooperative regime (i.e. $\mathbf{m}_{\rm int}=0$), the situation is completely different. Indeed when $\mathbf{m}_{\rm br}<\infty$ and the cooperation term $(b_{i})_{i\geq 1}$ has second moments, i.e.
$\bm{\sigma}^{2}_{{\rm int}}:=c+\sum_{i=1}^{\infty}i^{2}b_{i}<\infty$
we show that the associated BPI process does not come down from infinity. If $\mathbf{m}_{\rm br}<\infty$ and the cooperation term $(b_{i})_{i\geq 1}$ is a regularly varying sequence with index $\alpha\in(1,2)$ (i.e. $\bm{\sigma}^{2}_{{\rm int}}=\infty$), then we show that the speed of coming down from infinity is much faster than in the subcritical cooperative regime.
@@
Finally, in the subcritical cooperative case, we also show, under the assumption that the branching and cooperation parameters have large positive exponential moments, that the convergence of the speed of coming down from infinity also holds in the $\mathbb{L}_{p}$ sense for $p\in(0,2]$. Moreover, under the same hypothesis, we also deduce the second-order fluctuations around the speed of coming down from infinity for subcritical cooperative BPI processes at small times.
@@ 2 Main results
We first present our main results on the speed at which subcritical and critical cooperative BPI processes come down from infinity. In both cases, we characterise this rate under general assumptions. While the core techniques are fundamentally similar, key technical distinctions yield different speed rates and, consequently, require different parameter assumptions for the process to come down from infinity. In the second part of this section, we examine the second-order fluctuations in the speed of coming down from infinity within the subcritical cooperative regime, assuming the existence of large exponential moments on the branching and cooperation parameters.
@@ 2.1 The speed of coming down from infinity
Our first main result focuses on the subcritical cooperative regime, i.e. when $\mathbf{m}_{\rm int}<0$. Additionally, we assume that the branching parameter has a finite first moment, i.e., $\mathbf{m}_{\rm br}<\infty$. Under these assumptions, we have $c>0$, $\sum_{i\geq 1}ib_{i}<\infty$ and $\sum_{i\geq 1}i\pi_{i}<\infty$. Furthermore, the associated BPI process does not explode in finite time; see for instance Example 1 in [7] or Theorem 1 in [14].
@@
Examples of sequences that satisfy the previous hypothesis include regularly varying sequences. More precisely, let $\alpha\in(1,2]$ such that $b_{i}=i^{-\alpha-1}L(i)$, as $i\rightarrow\infty$, where $L:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is a slowly varying function at infinity. According to [9, Lemma 4.4], for any $\delta\in(1,\alpha)$,
$\sum_{i\geq n}i^{\delta}b_{i}\sim(\alpha-\delta)^{-1}n^{-(\alpha-\delta)}L(n),\qquad\textrm{as}\qquad n\rightarrow\infty.$
Using, for example, [9, (1) in Lemma 4.1] with $\varepsilon\in(0,\alpha-\delta)$, we find that $\sum_{i\geq n}i^{\delta}b_{i}\rightarrow 0$, as $n\rightarrow\infty$. This implies that $\sum_{i\geq 1}i^{\delta}b_{i}<\infty$, thus satisfying (${\bf H1}$) with $\alpha=\delta$.
@@
$\mathbb{P}\Big{(}Z^{(\infty)}_{t}<\infty,\,\,\,\,\text{ for all }\,\,t>0\Big{)}=1$
@@
$\lim_{t\downarrow 0}t\,Z_{t}^{(\infty)}=-\mathbf{m}_{\rm int},\qquad\text{almost surely}.$ (6)
@@
When branching and cooperation are absent (i.e. $d=\rho=0$ and $b=0$), Theorem 1 recovers the well-known result stating that the number of blocks in Kingman’s coalescent is asymptotically equivalent to $ct^{-1}$, almost surely as $t\downarrow 0$ (see, e.g., [6]). Recall that the BPI process corresponds to the block counting process of an EFC process with a coagulation mechanism featuring only the Kingman component and a fragmentation measure with no mass on partitions with infinitely many blocks, i.e. $c,\rho>0$ and $d=b=0$. Then, Theorem 1 provides the first result on the speed of coming down from infinity for EFC processes extending the known result from [5, Proposition 15] (see also [12, Corollary 1.2]), which asserts that the EFC process comes down from infinity.
In general, under the assumptions $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$ and (${\bf H1}$), Theorem 1 establishes that the Kingman component predominantly determines the speed at which the subcritical cooperative BPI process comes down from infinity. In contrast, the situation is markedly different for critical cooperative BPI processes, as discussed below. In the critical cooperative regime, the BPI process may fail to come down from infinity. Even in cases where it does, the speed differs significantly from that observed in the subcritical cooperative regime.
We now focus on the critical cooperative regime, i.e. we assume $\mathbf{m}_{\rm int}=0$ and $c>0$. It is important to note that the assumption $c>0$ is made to exclude the BGW process case, which does not come down from infinity. We also assume that the branching part of the BPI process satisfies
$\mathbf{m}_{\rm br}=-d+\sum_{i\geq 1}i\pi_{i}<\frac{\bm{\sigma}^{2}_{{\rm int}}}{2},$ (${\bf H2}$)
where $\bm{\sigma}^{2}_{{\rm int}}:=c+\sum_{i=1}^{\infty}i^{2}b_{i}\in(0,\infty]$. Note that (${\bf H2}$) implies that $\mathbf{m}_{\rm br}<\infty$. Hence, under (${\bf H2}$), the critical cooperative BPI-process does not explode in finite time; see [14, Theorem 1] or [7, Example 6].
Our second main result addresses two different scenarios: either $\bm{\sigma}^{2}_{{\rm int}}$ is finite or infinite. In the case where it is infinite, we specifically consider the regularly varying case, that is, there exists $\alpha\in(1,2)$ such that
$b_{i}=i^{-\alpha-1}L(i),\,\qquad\text{as}\,\qquad i\rightarrow\infty,$ (${\bf H3}$)
where $L:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is a slowly varying function at infinity. We observe that if (${\bf H3}$) holds, then necessarily $\bm{\sigma}^{2}_{{\rm int}}=\infty$ (one can use, for example, [9, Lemma 4.1 (1)]).
Theorem 2.
Suppose that $\mathbf{m}_{\rm int}=0,$ $c>0$ and that (${\bf H2}$) hold. Then, the family $(Z^{(z)})_{z\in\mathbb{N}_{0}}$ of BPI processes converges weakly, in $\mathbb{D}(\mathbb{R}_{+},\bar{\mathbb{R}})$, as $z\rightarrow\infty$, towards $Z^{(\infty)}=(Z^{(\infty)}_{t})_{t\geq 0}$, that is, a BPI-process started from $\infty$. Moreover,
  1. (i)
    If $\bm{\sigma}^{2}_{{\rm int}}$ is finite, then
    $\mathbb{P}\Big{(}Z^{(\infty)}_{t}=\infty,\,\,\text{ for all }\,\,t\geq 0\Big{)}=1.$ (7)
  2. (ii)
    If (${\bf H3}$) holds, then
    $\mathbb{P}\Big{(}Z^{(\infty)}_{t}<\infty,\,\,\,\text{ for all }\,\,\,t>0\Big{)}=1\qquad\text{and}\qquad\lim_{t\downarrow 0}\frac{Z_{t}^{(\infty)}}{v_{t}}=1,\quad\mathbb{P}-\text{almost surely}.$ (8)
    Moreover, for any $\varepsilon\in(0,2-\alpha)$, there are constants $t_{0},C_{1,\alpha,\varepsilon},C_{2,\alpha,\varepsilon}\in(0,\infty)$ such that
    $C_{1,\alpha,\varepsilon}t^{-\frac{1}{2-\alpha-\varepsilon}}\leq v_{t}\leq C_{2,\alpha,\varepsilon}t^{-\frac{1}{2-\alpha+\varepsilon}},\qquad\textrm{for}\qquad t\leq t_{0}.$
In particular, we have the following consequence of Theorem 2.
Corollary 1.
Suppose that $\mathbf{m}_{\rm int}=0,$ $c>0$, (${\bf H2}$) and (${\bf H3}$) hold. Then
$\lim_{t\downarrow 0}t\,Z_{t}^{(\infty)}=\infty,\qquad\mathbb{P}-\text{almost surely}.$ (9)
Corollary 1 shows that, under condition (${\bf H3}$), subcritical cooperative BPI-processes comes down from infinity faster than critical cooperative BPI-processes.

### v2

# On the speed of coming down from infinity for subcritical branching processes with pairwise interactions

@@ Abstract
In this paper, we study the phenomenon of coming down from infinity for subcritical cooperative branching processes with pairwise interactions (BPI processes) under suitable conditions. BPI processes are continuous-time Markov chains that extend classical branching models by incorporating additional mechanisms accounting for both competitive and cooperative interactions between pairs of individuals.
Our main focus is on characterising the speed at which BPI processes evolve when starting from a very large initial population in the subcritical regime. In addition, we investigate their second-order fluctuations. Furthermore, our results also apply to a class of exchangeable fragmentation-coalescent processes introduced by Berestycki [5] and several other models from population genetics.
@@ 1 Introduction
In this paper, we explore the speed of coming down from infinity for subcritical cooperative branching processes with pairwise interactions (referred to as BPI processes), under appropriate conditions. These processes were recently introduced in their current form by González-Casanova et al. [13] and further analysed by the authors in [7]. Our primary focus is on understanding their short-time behaviour when the initial population size is significantly large
BPI processes emerge as a natural extension of (continuous-time) Bienaymé-Galton-Watson process (BGW processes), addressing some of their inherent limitations, in particular their unrealistic properties. By incorporating pairwise interactions such as competition and cooperation between individuals, BPI processes can capture complex dynamics observed in biological systems. This generalisation has been studied in various contexts by several authors, including Jagers [19], Kalinkin [20, 21, 22, 23, 24], Lambert [27], and more recently by González-Casanova et al. [13] and the authors [7]. In these works, the birth and death rates are generalised by considering polynomial rates as functions of the population size which can be interpreted as different types of interactions between individuals.
Formally, let $\mathbb{N}\coloneqq\{1,2,\ldots\}$ and $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$. BPI processes are continuous-time Markov chains with state space $\mathbb{N}_{0}$. They feature natural births and deaths at rates proportional to the current population size, similar to the continuous-time BGW process, but also incorporate additional birth and death events due to cooperation and competition, occurring at rates proportional to the square of the population size. Specifically, a BPI process with parameters $c,d\in\mathbb{R}_{+}:=[0,\infty)$ and sequences $(b_{i})_{i\geq 1}$ and $(\pi_{i})_{i\geq 1}$ where $b_{i},\pi_{i}\in\mathbb{R}_{+}$, for $i\geq 1$, and $b\coloneqq\sum_{i\geq 1}b_{i}<\infty$ and $\rho\coloneqq\sum_{i\geq 1}\pi_{i}<\infty$, is defined by the infinitesimal generator $Q=(q_{i,j})_{i,j\in\mathbb{N}_{0}}$ given by
[…]
@@
From the infinitesimal generator (1.5), or $Q$-matrix, we observe that the additional death rate cz(z1)cz(z-1) represents competition, where one of the $z$ individuals selects another from the remaining z1z-1 at a constant rate $c$ and kills it. The additional birth rate bz(z1)bz(z-1) reflects cooperative interactions, where pairs of individuals produce $i$ new individuals with probability $b_{i}/b$.
@@
According to González-Casanova et al. [13], this quantity influences the long-term behaviour of BPI processes. A BPI process is termed supercritical, critical, or subcritical cooperative depending on whether $\mathbf{m}_{\rm int}>0$, $\mathbf{m}_{\rm int}=0$ or $\mathbf{m}_{\rm int}<0$, respectively. Theorem 1 in [13] shows that in the supercritical regime, and even under
$\mathbf{m}_{\rm br}\coloneqq-d+\sum_{i\geq 1}i\pi_{i}<\infty,$ (1.7)
@@
the BPI process may explode in finite time with positive probability. In the subcritical and critical cooperative regime, and under $\mathbf{m}_{\rm br}<\infty$, the BPI process does not explode in finite time almost surely. More recently, the authors [7] studied the explosion event for general subcritical and critical cooperative BPI processes (i.e. not necessarily satisfying that $\mathbf{m}_{\rm br}<\infty$) and provide general integral tests, see Theorems 1.1 and 1.4 in [7] for further details.
@@
When $c=b=0$, the BPI process reduces to the BGW process. A special case of BPI processes is the so-called logistic branching process (LB-process), deeply studied in [27], which corresponds to $c>0$ and $b=0$. Another example is the counting block process of the renowned Kingman coalescent, which corresponds to a BPI process with no branching or cooperation events (i.e., $c>0$ and $d=\rho=b=0$); see [25]. More generally, when $c,\rho>0$ and $d=b=0$, the BPI-process is exactly the block counting process of an exchangeable fragmentation-coalescence (EFC for short) process, introduced in [5], where the coagulation mechanism features only the Kingman component. Moreover, the BPI process can be viewed as the counting process of a particle system where particles can coagulate (via competition interactions), fragment without collision (via branching), or fragment due to collisions (via cooperative interactions). Ultimately, when $\pi_{i}=b_{i}=0$ for all $i\geq 2$, the BPI process reduces to a classic birth-and-death process.
@@
Motivated by studies on population dynamics with initially large populations, the phenomenon of coming down from infinity has attracted significant attention. This phenomenon describes a scenario where a random process starting from an infinite number of individuals (or blocks) immediately reaches a finite state. Kingman [25] was the first to observe this phenomenon in the context of the Kingman coalescent. Subsequently, Schweinsberg [33] provided necessary and sufficient conditions for the occurrence of this phenomenon in the setting of the $\Lambda$-coalescent. Further work, for the $\Lambda$-coalescent, includes studies of the speed of coming down from infinity by Berestycki et al. [6] and second-order asymptotics by Limic and Talarczyk [28, 29]. The case of exchangeable fragmentation-coalescence processes (EFC processes) was treated in Berestycki [5] where a sufficient condition for the process to come down from infinity is provided. Recent studies include those by Foucart [11], exploring the phenomenon of coming down from infinity for EFC processes with multiple coagulations but not simultaneous, as in the $\Lambda$-coalescent, and the fragmentation dislocates, at a finite rate, an individual block into sub-blocks of infinite size. Bansaye et al. [4] described the speed of coming-down from infinity for birth-and-death processes that eventually become extinct. More recently, Bansaye [3] developed a general framework for the phenomenon across various stochastic processes with jumps described by stochastic differential equations. Finally, it is important to note that the phenomenon of coming down from infinity for subcritical cooperative BPI processes was first studied by González-Casanova et al. [13] under the assumption that $\mathbf{m}_{\rm br}<\infty$. The authors [7] established a sufficient condition for general subcritical and critical cooperative BPI processes to come down from infinity. However, nothing is known about the speed at which BPI processes come down from infinity.
@@
Our aim is to determine the speed at which a subcritical BPI process comes down from infinity. In the subcritical cooperative regime (i.e. $\mathbf{m}_{\rm int}<0$), under the assumption that $\mathbf{m}_{\rm br}<\infty$ and the cooperation parameters $(b_{i})_{i\geq 1}$ have a bit more than a finite first moment (see ($\bf H1$})), we show that the speed of coming down from infinity is asymptotically equivalent to $\mathbf{m}_{\rm int}t^{-1}$. This is similar to the Kingman coalescent case where $\mathbf{m}_{\rm int}=c$.
@@
Finally, we also show, under the assumption that the branching and cooperation parameters have finite large positive exponential moments, the convergence in the $\mathbb{L}_{p}$ sense (for $p\in(0,2]$), of the subcritical cooperative BPI process rescaled by its speed of coming down from infinity. Moreover, we also obtain the second-order fluctuations around the speed of coming down from infinity at small times.
@@ 2 Main results
We first present our main results on the speed at which subcritical cooperative BPI processes come down from infinity. In the second part of this section, we examine the second-order fluctuations around this speed, assuming the existence of large exponential moments for the branching and cooperation parameters.
@@ 2.1 The speed of coming down from infinity
We henceforth consider the subcritical cooperative regime, i.e., when $\mathbf{m}_{\rm int}<0$. Additionally, we assume that the branching parameter has a finite first moment, i.e., $\mathbf{m}_{\rm br}<\infty$. Under these assumptions, we have $c>0$, $\sum_{i\geq 1}ib_{i}<\infty$ and $\sum_{i\geq 1}i\pi_{i}<\infty$. Furthermore, the BPI process does not explode in finite time; see for instance Example 2.1 in [7] or Theorem 1 in [13].
@@
Examples of sequences that satisfy the previous hypothesis include regularly varying sequences. More precisely, let $\beta\in(1,2]$ such that $b_{i}=i^{-\beta-1}L(i)$, as $i\rightarrow\infty$, where $L:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is a slowly varying function at infinity. According to Lemma 4.4 in [9], for any $\delta\in(1,\beta)$,
$\sum_{i\geq n}i^{\beta}b_{i}\sim(\beta-\delta)^{-1}n^{-(\beta-\delta)}L(n),\qquad\textrm{as}\qquad n\rightarrow\infty.$ (2.1)
By using, for example, (1) in Lemma 4.1 in [9] with $\varepsilon\in(0,\beta-\delta)$, we find that $\sum_{i\geq n}i^{\delta}b_{i}\rightarrow 0$, as $n\rightarrow\infty$. This implies that $\sum_{i\geq 1}i^{\delta}b_{i}<\infty$, thus satisfying ($\bf H1$}) with $\alpha=\delta$.
@@
$\mathbb{P}\Big{(}Z^{(\infty)}_{t}<\infty,\,\,\,\,\text{ for all }\,\,t>0\Big{)}=1$ (2.2)
@@
$\lim_{t\downarrow 0}(-\mathbf{m}_{\rm int})t\,Z_{t}^{(\infty)}=1,\qquad\text{almost surely}.$ (2.3)
@@
When branching and cooperation are absent (i.e., $d=\rho=0$ and $b=0$), Theorem 2.1 recovers the well-known result stating that the number of blocks in Kingman’s coalescent is asymptotically equivalent to $ct^{-1}$, almost surely as $t\downarrow 0$ (see, e.g., [6]). Recall that the BPI process corresponds to the block counting process of an EFC process with a coagulation mechanism featuring only the Kingman component and a fragmentation measure with no mass on partitions with infinitely many blocks, i.e. $c,\rho>0$ and $d=b=0$. Then, Theorem 2.1 provides the first result on the speed of coming down from infinity for EFC processes extending the known result in Proposition 15 of [5] (see also Corollary 1.2 in [11]), which asserts that the EFC process comes down from infinity. In general, under the assumptions $\mathbf{m}_{\rm int}<0$, $\mathbf{m}_{\rm br}<\infty$ and ($\bf H1$}), Theorem 2.1 establishes that the Kingman component predominantly determines the speed at which the subcritical cooperative BPI process comes down from infinity.
Our second main result in this section shows that the convergence (2.3) in Theorem 2.1 also holds in the $\mathbb{L}_{p}$ sense, for $p\in(0,2]$. However, this result requires the following additional assumption,
$\textrm{there exists}\quad\vartheta>e^{128\sqrt{2}}\quad\textrm{such that }\quad\sum_{i\geq 1}\vartheta^{i}\pi_{i}<\infty\quad\textrm{ and }\quad\sum_{i\geq 1}\vartheta^{i}b_{i}<\infty.$ ($\bf H4$})
It is important to emphasize that the previous assumption can be made for any value of $\vartheta>1$, though this introduces a dependency of $\vartheta$ on the values of $p$ for convergence within $\mathbb{L}_{p}$.
Theorem 2.2.
Suppose that $\mathbf{m}_{\rm int}<0$ and ($\bf H4$}) hold. Then, for $p\in(0,2]$, we have that
$\lim_{t\downarrow 0}\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}s\,Z_{s}^{(\infty)}-1\Big{|}^{p}\Big{]}=0.$ (2.4)
Indeed, the result in Theorem 2.2 holds for any $p>0$, provided that ($\bf H4$}) is satisfied for sufficiently large $\vartheta$ (see Section 3.3 for further details).